the mechanical power in an electric motor is given by P=T-I^2R. Find the current I(amperes) if the power P=448W, the total power T=480W, and the resistance R=2.00(Omega)
Answer by stanbon(75887)

the mechanical power in an electric motor is given by P=T-I^2R. Find the current I(amperes) if the power P=448W, the total power T=480W, and the resistance R=2.00(Omega)
---------------------------------
P = T-I^2R
---
448 = 480 - I^2*2
---
2*I^2 = 480-448
---
I^2 = 32/2
I^2 = 16
---
I = 4 amps
==============
Cheers,
Stan H.
============